OpenStudy (anonymous):

Can a 35 kg cannon ball and a 1.0 kg tennis ball have the same momentum? A) No, it is not possible. B) It is possible only is both have a constant acceleration in the y-direction. C) It is possible only if both balls are given the same initial velocity. D) It is possible if the tennis ball is given a larger initial velocity.

OpenStudy (mani_jha):

What is the definition of momentum? Try to answer this first

OpenStudy (unklerhaukus):

Momentum is mass times velocity \[P =mv\] if velocity is zero momentum is zero, this is independent of mass, hence the two objects have the same momentum when stationary

OpenStudy (ash2326):

Momentum is given as \[P=v\times v\] m= mass v= initial velocity Let m1 be the mass of cannon ball and its velocity be v1 Let m2 be the mass of tennis ball and its velocity be v2 We have to find a condition wherein both have the same momentum or \[m1 \times v1= m2 \times v2\] We are given m1= 35 kg and m2= 1 kg so \[35 \times v1=1 \times v2\] What do you think is the relation between v1 and v2? Which one should be larger?